# AI4EOSC: a Federated Cloud Platform for Artificial Intelligence in Scientific Research

### 17 Mar 2026 v2

@@ 4.1 AI4EOSC metadata
Once integrated, the Platform API retrieves the module’s metadata and serializes it in different formats (JSON-LD, RDF Turtle), being able also to transform it into different application profiles, such as the MLDCAT Application Profile [84]. This allows for a greater interoperability of our models with external catalogs, allowing semantic interoperability with formats that are compatible with MLDCAT-AP.
@@ 4.2.1 External deployment platforms
All executable assets in the platform are based on Docker containers, that are being released via reproducible builds into a public Docker Hub organization [^6] https://hub.docker.com/u/ai4oshub. This means that the AI modules (3.1.3) are easily deployable at external cloud providers and in any other infrastructure where Docker is available. From the Dashboard (3.1.2), we give users direct access to deploy in any external cloud provider via the Infrastructure Manager [32], that is part of our orchestration system (3.4). We have also developed connectors to be able to deploy on the EOSC EU Node [30], allowing users to leverage its resources.
@@ 4.2.2 External AI catalogs
Due to our AI modules design (3.1.3), which hides the module internal workings behind a unified API [37], we are able to easily build connectors with external AI module catalogs. These connectors allow to deploy those external modules in our platforms. To demonstrate this, we have built a connector to the BioImage Model Zoo catalog [66], that leverages their standardized metadata to map their functionality into our platform, and making them available directly in the Dashboard (3.1.2). This feature allows the allowing one-click deployments of those external modules, allowing the BioImage Model Zoo users to continue working with their tools, but making their outcomes available in the AI4EOSC marketplace.
@@ 4.2.3 External datasets
We implement this functionality transparently via sidecar tasks (3.1.1) and extending the the DataHugger tool [14] with additional integrations, currently including Zenodo [33], Data Europa [34], Dryad [28] or SeaNoe [82].
@@ 4.2.4 External storages
In addition to the platform storage, we are able to connect user deployments with any external storage, including Nextcloud, S3 [5], Owncloud [67], Dropbox [27], Google Drive [3], etc. In order to do so, we allow users to link the platform with the supported storage systems (e.g., NextCloud) or manually providing their user credentials for other storage providers.
@@ 5 Use cases and usage scenarios
5.1 Common usage scenarios
In this section we will explain how scientists can leverage the AI4EOSC Platform all along the ML lifecycle.
5.1.1 Developing a model
Developing an AI/ML model in AI4EOSC begins with a reproducible, MLOps-driven workflow that ensures traceability and compliance with FAIR principles. Researchers and platform users can deploy a pre-configured development environment (Codespace) via the Dashboard, selecting their preferred IDE (VSCode or JupyterLab) and utilizing their preferred framework (e.g., TensorFlow, PyTorch). The environment is linked to version-controlled repositories (GitHub) and experiment tracking (MLflow) tools, enabling users to capture metadata into it. The repository structure is set up based on one of the AI4EOSC software templates.
As already mentioned, the platform enforces a structured JSON schema for model metadata (e.g., dependencies, datasets, licenses), ensuring interoperability with external catalogs (MLDCAT-AP). Every code commit triggers the CI/CD pipeline, that ensures that the development steps are logged into the W3C PROV-compliant provenance graph, enabling full lineage tracking. Moreover the same pipeline, enforced at platform level, builds the Docker image and refreshes all the relevant components, so that the new module is ready to be used.
5.1.2 Training a model
Training in AI4EOSC leverages our federated Workload Management System (WMS) to dynamically allocate resources (GPUs and CPUs) across distributed providers, resource allocation, energy consumption and the overall performance. Researchers configure training jobs via the Dashboard, specifying the data sources that need to be linked with their deployment (e.g., Zenodo datasets), using data annotation tools within the platform, or utilizing federated learning by leveraging the integrated federated learning server.
The generated trained model can then be integrated in the module catalog, as in the model development case (5.1.1), performing the same MLOps and CI/CD approach previously described.
5.1.3 Deploying a module
A deployment in AI4EOSC is provided via different options, according to the user and application needs, and the resources involved. AI4EOSC covers the computing continuum in a multi-faceted approach:
  • 1.
    Permanent deployment within the WMS, if they want their module to be permanently deployed for low-latency or prototyping.
  • 2.
    External platforms with one-click deployment to external or commercial clouds (AWS, GCP) via the Infrastructure Manager [16]
  • 3.
    On-demand inference using the AI as a Service component (3.2) based on OSCAR, following a serverless MLOps-optimized approach with on-demand autoscaling.
  • 4.
    Edge/Cloud Continuum: Models can be deployed to HPC clusters or edge devices (Raspberry Pi [75]).
@@ 5.2 Platform adoption
Besides building a generic and horizontal platform, and in order to foster the adoption of the technology, we have closely worked with several initiatives in order to build a customized version the platform, tailored to the user community needs. The interactions with those communities has led to a virtuous circle, where the platform has guided them on their adoption of AI and they have in turn provided feedback for a user-driven platform development.
@@
The iMagine [^7] https://www.imagine-ai.eu/ initiative is a European Union–funded research project that focuses on developing and operating an AI-enabled imaging platform for aquatic science. It aims to give researchers working on oceans, seas, coastal and inland waters access to rich image datasets, advanced AI-powered analysis tools and best-practice guidance for scientific image processing. The iMagine platform is built on the software stack developed within AI4EOSC project.
During iMagine, the Competence Center team provided support to the usersa and collaborated closely with the developers of AI4EOSC project, providing feedback to enhance the platform. This partnership enabled hands-on experience across the full AI workflow, from image collection and dataset preparation to model training and deployment for aquatic science applications, ultimately resulting in the publication [8].
@@ 5.2.2 KMD4EOSC
KMD4EOSC [^8] https://kmd4eosc.pl/ is a Poland national project for data storage and sharing, and efficient processing of large volumes of data in HPC, BigData, and AI Models. The project is co-funded by European Funds for a Modern Economy Program and the European Union. The project is one of the flagship projects of the Polish Research Roadmap.

### 05 May 2026 v3

@@ 4.1.1 AI4EOSC metadata
Once integrated, the Platform API retrieves the module’s metadata and serializes it in different formats (JSON-LD, RDF Turtle), being able also to transform it into different application profiles, such as the MLDCAT Application Profile. This allows for a greater interoperability of our models with external catalogs, allowing semantic interoperability with formats that are compatible with MLDCAT-AP.
4.1.2 Provenance
The system is designed to operate in a non-intrusive manner, leveraging metadata already generated by existing architectural components such as MLflow and Jenkins, without requiring modifications to these systems. Once a module has committed changes to the Git repository, the CI/CD pipeline triggers the collection of provenance metadata from multiple sources: module metadata (4.1.1) and training resources (3.1.1) from the Platform API (3.1.2), experiment tracking metrics from MLflow (3.1.5), and build information from the CI/CD pipeline itself (3.1.5).
These heterogeneous, non-semantic metadata fragments are stored in their native JSON representation in a PostgreSQL database. This design choice is deliberate: rather than imposing a rigid schema at ingestion time, the system preserves the original structure of each source, enabling flexible querying and transformation. Semantic interoperability is then achieved by applying RDF Mapping Language (RML) rules via CARML over the stored JSON fragments. These declarative mappings define how disparate metadata elements correspond to a unified semantic model, enabling the integration and linkage of otherwise disconnected provenance information into a coherent end-to-end provenance graph.
The resulting provenance graph is serialized as an RDF document encoded in JSON-LD, ensuring compliance with the PROV-O ontology for standardized provenance modeling. The system also supports alignment with domain-specific ontologies, allowing for contextual enrichment and adaptability across different AI application domains. Having the provenance data in RDF further allows conversion to external formats such as FAIR4ML [16] or any additional mapping, and can be extended with virtually any information that the platform or the user is able to provide.
A comparative analysis of this design against other PROV-based ML pipeline approaches is presented in Section 2.
In order to present the provenance information to users, an interactive web-based visualization tool has been developed and integrated into the AI4EOSC dashboard, providing dynamic, user-friendly interfaces for navigating complex provenance graphs. Additionally, users can query the provenance graph in natural language via integration with the chatbot assistant (Section 3.3), enhancing transparency, traceability, and reproducibility within AI deployment workflows.
Figure 9: Example of a provenance graph built for a demo application.
Figure 9 traces a concrete provenance record for the exemplary ai4os-demo-app module. When a developer commits new code, the Jenkins CI/CD agent detects the change (jenkins-polling), triggering Build 101. This activity, that is associated with Jenkins acting on behalf of the AI4EOSC platform, validates metadata, builds a Docker image, and notifies the provenance system. Simultaneously, a training run (UUID 5e96fe94) executes on Nomad resources with metrics logged to MLflow, producing a versioned build artifact. All these relationships are captured as W3C PROV predicates (e.g. wasGeneratedBy, wasAssociatedWith, etc.), forming a complete, queryable lineage graph for the module.
4.1.3 Openness
AI4EOSC integrates open-source principles at its core: all the platform software is publicly available [^5] https://github.com/ai4os, which combined with the provided TOSCA deployment recipes used in the PaaS (3.4), enables communities to redeploy and built on top of our platform.
In a similar spirit, every AI module built with the platform has a public code repository attached to it, hosted under a common GitHub organization [^6] https://github.com/ai4os-hub. The code, as as well as any other downloadable asset (model weights, training dataset, metadata, provenance file, etc) is clearly linked from the AI module page in the AI4EOSC Dashboard (3.1.2). To avoid vendor lock-in, AI4EOSC only allows but facilitates the redeployment of all AI modules in external clouds (4.2.1).
@@ 4.2.1 External deployment platforms
All executable assets in the platform are based on Docker containers, that are being continuously released via reproducible builds into a public Docker Hub organization (3.1.5). This means that the AI modules (3.1.3) are easily deployable at external cloud providers and in any other infrastructure where Docker is available. From the Dashboard (3.1.2), we give users direct access to deploy in any external cloud provider via the Infrastructure Manager, that is part of our orchestration system (3.4). We have also developed connectors to be able to deploy on the EOSC EU Node, allowing users to leverage its resources.
@@ 4.2.2 External AI catalogs
Due to our AI modules design (3.1.3), which hides the module internal workings behind a unified API [28], we are able to easily build connectors with external AI module catalogs. These connectors allow to deploy those external modules in our platform. To demonstrate this, we have built a connector to the BioImage Model Zoo catalog [41], that leverages their standardized metadata to map their functionality into our platform, and making their AI modules directly available in our Dashboard (3.1.2). This feature allows one-click deployments of those external modules, allowing the BioImage Model Zoo users to continue working with their tools, but making their outcomes available in the AI4EOSC marketplace.
@@ 4.2.3 External datasets
We implement this functionality transparently via WMS sidecar tasks (3.1.1) and extending the the DataHugger tool [12] with additional integrations, currently including Zenodo, Data Europa, Dryad or SeaNoe.
@@ 4.2.4 External storages
In addition to the platform storage, we are able to connect user deployments with any RCLONE-compatible external storage, including Nextcloud, S3, Owncloud, Dropbox, Google Drive, etc. In order to do so, we allow users to easily link the platform with selected storage systems (e.g., NextCloud) or manually provide their user credentials for other storage providers.
4.3 Security and access control
Taking into account the multi-user, multi-institution and multi-country nature of the AI4EOSC platform, security and data protection have been carefully considered in the design and implementation. In this regard, the platform adopts a layered security model that addresses authentication, authorization, secret management, tenant isolation, and data protection.
4.3.1 Authentication and authorization
The platform enforces authentication and authorization through dedicated Keycloak instance, using OpenID Connect (OIDC) as protocol. This way, all AI4EOSC services use this endpoint as the authoritative source of authentication, acting as a single trusted broker with external federated identity providers, such as the EGI Check-In or the MyAccessID (eduGAIN) services. This design, aligned with the AARC Blueprint Architecture [1], reduces the trust boundary to a single managed service, reducing the complexity of dealing with identity providers at the individual service level, thus reducing the exposure level.
Authorization follows a role based access control (RBAC) derived both from the internal group mapping in the system and from the inherited entitlements from trusted identity providers. The platform services enforce different tiered access levels (from short-lived and constrained demo access to complete resource access).
4.3.2 Secret management
Besides user authentication, the platform manages a set of other secrets that can grant access to different components of the platform. These include storage tokens used to sync with external providers or tokens to join a federated learning scenario, among others. In order to manage them security, AI4EOSC enforces the usage of Vault, so that secrets are never exposed as plaintext to the platform infrastructure, as they are injected via the WMS (3.1.1).
AI4EOSC does not enforce credential rotation, but users can define the secret lifetime when creating them and secrets can be revoked at any time via the platform dashboard or the API. This allows, for instance, to invalidate a token of a malicious client participating in a federated learning scenario or to revoke a compromised credential. This authentication is also intended to prevent poisoning attacks by unauthorized clients attempting to connect to the training, as will be further explored in the Section 4.4.3.
4.3.3 Tenant isolation
AI4EOSC provides two levels of isolation. At the organizational level, Nomad namespaces are leveraged to partition resources across dedicated platform instances, for example for different research groups, institutions, or projects such as iMagine or KMD4EOSC (6.2). This ensures that the resources, catalogs, and configurations of different communities are fully segregated. Moreover, at workload level, user deployments are isolated at the internal network level, ensuring that deployments from different users do not interfere with, or observe, each other. Network-level access between deployments is restricted by default, so that inter-deployment communication is only possible through the explicitly exposed public endpoints, protected by the token-based access control described before.
4.4 Privacy considerations
For a scientific platform like AI4EOSC, privacy is a paramount concern, as researchers frequently handle sensitive, proprietary, or geographically restricted datasets (e.g. medical datasets). Ensuring robust privacy mechanisms not only guarantees compliance with European data protection regulations such as the GDPR, but also establishes the necessary trust among institutions to enable collaborative AI workflows without compromising data confidentiality.
4.4.1 Data retention
According to the GDPR, the platform has been designed following the data minimization platform. Uploaded data is only stored in the user personal space and it is not retained after the user deletes it. When a user deployment is terminated, all locally cached data in the compute nodes (including datasets synchronized from external data providers) is completely removed. Users are required to adhere to the AI4EOSC privacy policy upon registration, and data processing agreements are in place with all federated infrastructure providers.
4.4.2 Logging and traceability
All user-facing actions are always routed through the Platform API (PAPI) (Section 3.1.2), which provides a centralized and structured audit log of all platform activity. This includes job submissions, deployment lifecycle events, secret creation and revocation, and catalog interactions, enabling operators to trace the origin of any platform action.
4.4.3 Federated learning
Federated learning is a privacy-preserving approach by design, as it does not require participating clients to share their data with each other or with a central server. However, additional security and privacy risks remain, including poisoning attacks and client-side inference attacks, which must be addressed at the platform level.
As described in Section 3.1.4, AI4EOSC extends Flower with token-based authentication and client weight divergence monitoring to defend against poisoning from both unauthorized and malicious authenticated clients. Server-side differential privacy and metric privacy [48] further protect against inference attacks on the aggregated model. Together, these mechanisms form a layered defense that operates transparently within the platform without requiring additional configuration from researchers. More specific measures remain complex to implement generically at the platform level and need to be addressed on a case-by-case basis.
4.5 Reliability, Fault Tolerance, and Elasticity
The distributed and federated nature of the AI4EOSC platform requires a robust and reliable resource management system. Our design prioritizes system graceful degradations to maintain service availability despite the heterogeneous reliability of the underlying higly distributed provider nodes.
4.5.1 Fault Tolerance and Self-healing
To address potential failure of nodes in the WMS, we leverage the native self-healing capabilities of the underlying orchestrator (i.e., HashiCorp Nomad, as described in Section 3.1.1). The system maintains a constant state and heart-beat monitor of all federated nodes and supports graceful degradation due to the Nomad server redundancy. Additionally, Nomad incorporates consensus protocols to elect new cluster leaders in the event of a failure or network partitions.
4.5.2 Resource Elasticity and Backpressure
The platform differentiates between the elasticity requirements of training and inference workloads:
  • 1.
    AIaaS Autoscaling: For inference services, the platform implements reactive horizontal autoscaling. Based on real-time metrics (e.g., request latency or CPU/GPU utilization), the AIaaS layer can dynamically provision additional container instances to handle traffic spikes, ensuring low-latency responses for end-user applications.
  • 2.
    WMS Queueing and Determinism: For training workloads, we adopt a deterministic resource allocation strategy. Given the high cost and limited availability of GPU resources across the scientific resource providers, we do not implement automated hardware provisioning (cluster autoscaling). Instead, we manage high demand through a queueing system. This prevents the over-saturation of the system as new jobs are held in a pending state until sufficient resources are released by completed tasks.
5 Performance evaluation
5.1 Scalability
To asses the scalability of the platform, in Figure 10 we show the time it takes to search and retrieve the full information of a single job, as a function of the number of the total jobs deployed in the AI4EOSC cluster federation. In blue we show the scenario when all the jobs are running in the same datacenter, while in red we show the case where the jobs are equally distributed across all the datacenters of the federation. As can be seen, search times increase sub-linearly with the number of total jobs, while job retrieval times stay constant, both in the range of few tens of milliseconds. More importantly, increasing the number of datacenters in the federation only increases search times by a small fixed cost. This, together with the sublinear increase in search times, shows the robust scalability capacity of the AI4EOSC federation, both in the number of jobs managed, as well as in the number of integrated infrastructure providers.
Figure 10: Search and retrieval times for a single WMS deployment, as function of the total jobs in the cluster. In blue, when launching all the jobs in the same datacenter. In red, when distributing the load equally among all datacenters. For clarity, a rolling mean average with window size 10 has been applied to smoothen the curves.
5.2 Training Efficiency
To test the possible overhead of training in an AI4EOSC environment, we train the same Computer Vision model (from a real-life AI4EOSC usecase published in [49]) in three different scenarios: (1) directly in a Virtual Machine (VM); (2) inside a Docker container deployed in a Virtual Machine; and (3) in an AI4EOSC Development Environment (3.1.4), which is, in the end, a Docker container running in a VM with Nomad/Consul orchestration (3.1.1) and accessed via VScode (3.1.4).
In each scenario, the model is trained for 25 epochs on the same resources (one NVIDIA T4 16GB GPU, 20 CPU cores, 86 GB RAM, 40 GB disk) with the same environments (Python 3.13, TensorFlow 2.20, CUDA 12.4, cuDNN version 9.21.01). The dataset is roughly 1 GB in size.
VM VM + Docker AI4EOSC
$12.4\pm 1.7 11$$2.6\pm 1.2 11$$13.5\pm 1.2$
Table 2: Training time (in seconds) for a real life AI4EOSC usecase trained for 25 epochs, averaged over 5 training runs.
As shown in Table 2, the dockerization on top of the VM does not add a significant difference in training times. Running in an AI4EOSC environment does result in a minor time increase (less than 1%). This very small difference, caused by the Nomad/Consul orchestration overhead, demonstrates that the AI4EOSC platform is able to transparently serve federated resources at scale to users without significant performance loss.
5.3 Federated Learning Efficiency
Concerning the Flower FL server and the efficiency of the implementation in AI4EOSC, we have conducted four experiments with the same use case, considering in this set up the four clients presented in [49]. This experiments are as follows: (1) both the FL server and the four clients are deployed in AI4EOSC; (2) the FL server is deployed in AI4EOSC and the four clients run in four VMs; (3) both the FL server and the four clients run in five different VMs; (4) the FL server is deployed in a VM and the clients run in four separated AI4EOSC deployments. Concerning the training, each client trains the model for one epoch, and the FL process is conducted during five rounds. The results obtained in the four setups (mean and standard deviation after repeating the experiments five times) are shown in Table 3.
Server Clients AI4EOSC VMs
AI4EOSC 229.18 $\pm$ 5.90 233.60 $\pm$ 14.61
VM 222.61 $\pm$ 14.16 226.96 $\pm$ 11.49
Table 3: Training time (in seconds) for a federated learning setting with 5 rounds, averaged over 5 training runs with the standard deviation.
Regarding the client-side resources used in these scenarios, in AI4EOSC we deployed four instances, each configured with 4 vCPUS, 10 GB of RAM and 20 GB of disk memory. These specifications were chosen to ensure comparability with the selected flavor in the four VMs, matching the number of VCPS and disk capacity, and providing a similar RAM memory (10.7GB). In all the cases the same version of the FL server and the dependencies for running the clients were used (including Flower extensions).
[…]
Finally, these results indicate that using the AI4EOSC platform in this FL setting does not introduce a significant computational overhead, with training times remaining comparable across configurations.
5.4 Cold-start
Cold-start latency is observed mainly in the first invocations after service deployment or scale-up, when the platform must complete additional preparation steps such as image handling, container startup, and initialization before the request can be served. Cold-start is particularly important for the AIaaS serverless component (based on OSCAR), as it handles short-lived inference calls (spawn container, make an inference, kill the container). While execution times are application and container image-dependent, for a reference implementation, synchronous parallel OSCAR invocations complete with sub-second latency, indicating that the platform can sustain interactive request-response workloads when the service is already warm and resources are available. This was achieved in a 3-node OSCAR cluster (24 cores and 54 GBs of RAM available), showing an average first synchronous invocation time of 1521 ms and warm synchronous invocations of 992 ms.
Cold-start latency is not critical in the WMS, because the WMS is meant to executed longer jobs and tasks, like AI trainings, and thus user experience is not meaningfully impacted by warm/cold start delays (typically in the few seconds if the image is pre-cached in the node, to tens of seconds if the full image has to be pulled from the AI4EOSC container registry).
@@ 6 Use cases and usage scenarios
6.1 End-to-end user journey: precipitation nowcasting with federated learning
To better illustrate how a researcher can interact with the AI4EOSC platform across the full AI/ML lifecycle, in this section we describe a concrete use case focused on radar-based precipitation nowcasting [49]. This use case aims to represent a broad class of scientific scenarios in which training data is distributed across institutions and cannot be centralized due to technical, legal, or administrative constraints.
Problem and data
The goal of the selected case is to predict precipitation, expressed as Vertically Integrated Liquid (VIL), in a five minute horizon, using sequences of radar images captured by a meteorological radar located in the borderland between the Czech and Slovak Republic. The dataset covers four months of radar observations (between April and July 2016), with images captured every five minutes and stored as HDF5 files. To simulate a realistic federated scenario, the images are divided into four quadrants, each treated as a separate client with a distinct and heterogeneous data distribution. The goal of the case is to validate the feasibility of using federated learning for such predictions, where data is not directly accessible.
Data was stored in the storage system of the platform, that can be securely accessed from the user training jobs. In a realistic scenario, users would have been able to link their external storage system with their accounts, so that their datasets were not centrally stored, but rather accessed via their own storage systems only at training time. As outlined in Section 4.4, data is not retained after the training process is finished.
Development
The researchers deployed a Codespace via the AI4EOSC Dashboard, selecting a TensorFlow environment running on an NVIDIA Tesla T4 GPU (16 GB RAM, 8 CPU cores, 10 GB disk). The development environment provided secure and authenticated access to version-controlled repositories and the MLflow experiment tracking service, allowing the team to log and compare training runs across different model configurations and hyperparameter settings.
Federated training
Once a suitable convolutional neural network architecture was identified and validated, the team configured a Flower federated learning server via the Dashboard, deploying four client instances within the platform itself, one corresponding to each zone. Token-based authentication was used to ensure that only authorized clients could participate in the training. The FL process ran for ten rounds of ten epochs each. Following the federated phase, a novel personalized FL approach (adapFL) was applied, further fine-tuning the aggregated global model locally on each zone data for a set of ten additional epochs. This adaptive step allowed each client to recover zone-specific characteristics lost during federated aggregation.
Deployment and results
In order to quickly validate the results, the generated models were deployed as permanent WMS endpoints for prototyping and evaluation, and as serverless AIaaS endpoints via OSCAR for on-demand inference. The platform CI/CD pipeline automatically built and published the corresponding Docker images, validated the module metadata, and triggered provenance graph updates, ensuring full traceability of the training process without any additional intervention from the researchers. The adapFL approach yielded improvements in terms of both metrics evaluated (MSE and MAE) over individual training and vanilla FL in all four zones, demonstrating that privacy-preserving collaborative training on heterogeneous distributed data can yield better predictive models than local training alone.
@@ 6.2 Platform adoption
Besides building a generic and horizontal platform, and in order to foster the adoption of the technology developed within multidisciplinary domains, we have closely worked with several initiatives. The aim is for these communities to collaborate on the co-design of the platform in order to build a customized version that adapts to the needs of the user community. The interactions with those communities has led to a virtuous circle, where the platform has guided them on their adoption of AI and they have in turn provided feedback for a user-driven platform development.
@@
The iMagine [^7] https://www.imagine-ai.eu/ initiative is an European Union–funded research project that focuses on developing and operating an AI-enabled imaging platform for aquatic science. It aims to give researchers involved in the study of oceans, seas, coastal and inland waters access to rich image datasets, advanced AI-powered analysis tools and best-practice guidance for scientific image processing. The iMagine platform is built on top of the software stack developed within the AI4EOSC project.
Throughout the run of the iMagine project (which formally concluded in August 2025), the Competence Center team provided support to the users and collaborated closely with the developers of the AI4EOSC project, providing feedback to enhance the platform. This partnership enabled hands-on experience across the full AI workflow, from image collection and dataset preparation to model training and deployment for aquatic science applications, ultimately resulting in the publication [7]. It is important to note that the project has developed several use cases built on top of the services provided within the platform, including [29] (focusing on satellite-derived bathymetry mapping) and [24] (focusing on distance classification of marine vessels) among others.
@@ 6.2.2 KMD4EOSC
KMD4EOSC [^8] https://kmd4eosc.pl/ is a Polish national project for data storage and sharing, and efficient processing of large volumes of data in HPC, BigData, and AI Models. The project is co-funded by European Funds for a Modern Economy Program and the European Union. The project is one of the flagship projects of the Polish Research Roadmap.